Peach Corporation is considering several investment proposals, as shown below: Investment Proposal A B C D Investment required $ 132,000 $ 165,000 $ 99,000 $ 123,750 Present value of future net cash flows $ 158,400 $ 247,500 $ 138,600 $ 276,000 If the project profitability index is used, the ranking of the projects from most to least profitable would be:
Answer #1

Project profitability index = Net present value of future cash inflows / Investment required

A= 158,400/132,000 = 1.2

B= 247,500/165,000 = 1.5

C= 138,600/99,000 = 1.4

D= 276,000/123,750 = 2.2

A, C, B, D

Comment if you face any issues